OpenStudy (anonymous):

help please? Jacob sells homemade skateboard decks for a profit of $27 per deck. He is considering switching to a new type of material that will increase his profit, as expressed by the function y = 32x, where x is the number sold and y is the amount of profit. How many more dollars will Jacob earn on each skateboard deck if he switches to the new material? Explain how you found your answer.

OpenStudy (danjs):

5 per board

OpenStudy (danjs):

the amount earned per board is the slope of the line, which in each case is 27 and 32

OpenStudy (danjs):

y = 27*x and y = 32*x

OpenStudy (danjs):

Put in x=1 board sold, how much more is the profit in the second case than the first case?